OpenStudy (anonymous):

I'm reading my lesson over because I forgot how to go about some stuff and I come across this, is this technically considered wrong?

OpenStudy (anonymous):

Technically 0 squared is 0, a square root cannot be 0. It's any whole number. 1,7,9 etc. Even though they might be irrational.

OpenStudy (agent0smith):

0 is an integer. 0^2 = 0 and the square root of zero is zero.

OpenStudy (anonymous):

Yeah, but a square root cannot be 0, its any positive number.

OpenStudy (agent0smith):

The square root of zero is zero, so... why^^?

OpenStudy (anonymous):

0 is not a positive number, its in between negative and positive, neutral.

OpenStudy (agent0smith):

Why does a square root have to be a positive number, as opposed to a positive number or zero?

OpenStudy (anonymous):

Because a square root cannot be negative. I don't consider 0 to be positive. It's an integer, yes, 0^2 = 0 and the sqrt of 0 is 0 but idk :l

OpenStudy (agent0smith):

0 isn't negative. A square root can be positive, or zero. I don't see an issue at all.

OpenStudy (anonymous):

I don't know how 0 can be a square root lol, but I'll leave it at that. Thanks for answering my doubt (:

OpenStudy (agent0smith):

I don't know how it CAN'T be. 0*0 = 0 Sqrt(0) = 0

OpenStudy (agent0smith):

Unless you're talking about the limit as x approaches zero, since you can only approach from the right, not the left... i don't see the problem.

OpenStudy (anonymous):

:D I guess tehe ty

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=square+root+definition&a=*MC.~-_*MathWorld- I still can't figure out why you think it can't be zero :P